OpenStudy (anonymous):

what is the interquartile range of the following data set? 46, 9, 48, 96, 61, 84, 29, 1, 82, 5, 34

OpenStudy (ajprincess):

First arrange these data in ascending order. Then find Q1 and Q3. Q1=(N+1)/4 Here Q1-first quartile and N-total number of data. Q3=3/4*(N+1) Here Q3-third quartile Interquartile range = Q3-Q1.